HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

UNAMBIGUOUS ASSIGNMENT OF ETHYLENE TRANSITIONS IN THE 5900-6200 CM^{−1} RANGE

\underline{S. PEROT} , J. LECOMTE, N. SUAS-DAVID, L. RUTKOWSKI, R. GEORGES, IPR, UMR 6251, Université de Rennes, Rennes, France; M. REY, GSMA, UMR 7331, Université de Reims, Reims, France; V. TYUTEREV, V.E. Zuev Institute of Atmospheric Optics SB RAS, Tomsk, 634055, Russia; S. KASSI, LIPhy, UM R5588,

Université Grenoble Alpes, Saint Martin d’Hères, France; V. BOUDON, C. RICHARD, ICB, UMR 6303, Université de Bourgogne, Dijon, France

Among other species, methane^{1} and acetylene^{2} were detected in the atmospheres of Hot Jupiter exoplanets. The James Webb Space Telescope is expected to detect other small hydrocarbons including ethylene (C_{2}H_{4}), in particular in the near-infrared spectral range^{3}. High-temperature and high-resolution laboratory data are therefore needed both to unambiguously detect this species and to retrieve its atmospheric concentration and temperature profiles. Our aim is to produce high-temperature spectroscopic data of ethylene in the 1.6-1.7 μm spectral region using our SMAUG hypersonic wind tunnel, which has already been applied to methane^{4}^{,}^{5}. Nevertheless, due to the very dense absorption spectrum of ethylene even at room temperature, a preliminary measurement campaign was carried out at very low temperatures under slit jet-cooled conditions to first provide accurate assignments of cold band transitions. A series of three spectra were recorded by cavity ringdown spectroscopy and tunable distributed feedback laser diodes at rotational temperatures of 6 K, 16 K and 54 K using either Ar/C_{2}H_{4} or N_{2}/C_{2}H_{4} gas mixtures. Unambiguous assignment of the observed rovibrational lines is performed using the TheoReTS ab initio line list^{6} and the SPVIEW/XTDS software package^{7}. These transitions have been attributed to seven interacting vibrational bands: ν_{2} +ν_{3} +ν_{11}, ν_{2} +ν_{6} +ν_{9}, ν_{1} +ν_{11}, ν_{5} +ν_{11}, ν_{1} + ν_{2} + ν_{12}, ν_{5} + ν_{9}, and ν_{9} + 2ν_{12}.

Support from the ANR-RNF TEMMEX project (grant ANR-21-30CE-0053-01 and grant RSF 22-42-09022) is acknowledged.

1doi:10.1038/nature06823, M. Swain, G. Vasisht, G. Tinetti, Nature, 452, 329-331 (2008).

2doi:10.1038/s41586-021-03381-x, P. Giacobbe et al., Nature, 592, 205-208 (2021).

3doi:10.1051/0004-6361/202141468, D. Gasman,et al., A&A, 659, A114 (2022).

4doi:10.1063/5.0003886, E. Dudás et al., J. Chem. Phys., 152, 134201 (2020).

5doi:10.1016/j.icarus.2022.115421, E. Dudás et al., Icarus, 394, 115421 (2023).

6doi:10.1016/j.jms.2016.04.006, M. Rey, et al., J. Mol. Spectrosc., 327, 138-158 (2016)

7doi:10.1016/j.jms.2008.01.011, Ch. Wenger, et al., J. Mol. Spectrosc., 251, 102-113 (2008)

p-number: p120  Submitted on Wed, 14 Jun 2023 10:18:15 +0200

p-number: p120  Submitted on Wed, 14 Jun 2023 10:18:15 +0200